OpenStudy (anonymous):

-16x^4-32x^3-80^2 factoring help asap please

OpenStudy (anonymous):

What common factors do you see?

OpenStudy (anonymous):

Let's not just toss out the answer, or we'll have ten more problems like this to do.

OpenStudy (anonymous):

i dont even know how that person did that, i need explanation

OpenStudy (anonymous):

oh shoot i meant 80x^2

OpenStudy (texaschic101):

you need to find a common factor that will go into each of the numbers

OpenStudy (anonymous):

8x?

OpenStudy (anonymous):

i wasnt tryna be rude or start anything i was js

OpenStudy (anonymous):

nooooo wait16x?

OpenStudy (texaschic101):

try -16x^2

OpenStudy (anonymous):

You can always divide out a factor and see whether there still is one. What trinomial do you get after removing the common factors?

OpenStudy (anonymous):

grr, your right,

OpenStudy (anonymous):

x^2-16x-5?

OpenStudy (anonymous):

i mean x^2+2x-5

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

oh ok, ya the first one confused me but yah

OpenStudy (anonymous):

yah

OpenStudy (anonymous):

nope u lost me, the drawing confused me

OpenStudy (anonymous):

May I step in?

OpenStudy (anonymous):

sure

OpenStudy (anonymous):

\[-16x^4-32x^3-80x^2=-16(x^2+2x^3+5)\] Multiply it out to test it. Then tell me if you follow so far.

OpenStudy (anonymous):

i follow

OpenStudy (anonymous):

i give up

OpenStudy (anonymous):

well i really need to know this its just confusing with different people telling me different stuff

OpenStudy (anonymous):

I left out the x^2. Let me fix it.

OpenStudy (anonymous):

\[-16x^4-32x^3-80x^2=-16x^2(x^2+2x+5)\]

OpenStudy (anonymous):

@numb3r1 i follow now

OpenStudy (anonymous):

@Loser66 i feel like you are blowing things out of proportion thank you for ur help

OpenStudy (anonymous):

OK, so keep the 16x^2 off to the side. Can the trinomial be factored, either manually or with the quadratic formula?